OpenStudy (lexy148):

I need help please (this is due by tommorrow) A bag contains five red chips, three white chips, and two blue chips. Three chips are to be selected at random, without replacement. Find the probability that a. all are red b. the first two are red and the third is blue c. the first is red, the second is white and the third is blue

OpenStudy (anonymous):

this is not nearly as hard is it seems, as long as you know how to compute 10 choose 3 and stuff like that you know how to do that?

OpenStudy (lexy148):

um no i dont

OpenStudy (anonymous):

oh nvm now that i actually read the question, you don't need that at all

OpenStudy (anonymous):

lets do the first one what is the probability that the first one selected is red?

OpenStudy (lexy148):

5/10 wich is reduced to 1/2

OpenStudy (anonymous):

ok now we have a red one in our hand (because we are not replacing it) what is the probability that the next one chosen is also red?

OpenStudy (lexy148):

4/10

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

there are 4 red ones left for sure but how many chips are left in the bag?

OpenStudy (lexy148):

thers 9 left in the bag

OpenStudy (anonymous):

zactly would you like to change your answer from \(\frac{4}{10}\) ?

OpenStudy (lexy148):

if thats the correct answer i would change it

OpenStudy (anonymous):

no 4/10 was incorrect since there are 4 red ones left out of the total of 9, the probability that the next one is red is \(\frac{4}{9}\)

OpenStudy (anonymous):

one more to go we took out 2 red chips how many red chips are left, how many chips total?

OpenStudy (lexy148):

3 left, 7 total

OpenStudy (anonymous):

i think 8 total since you started with 10 have have taken out two

OpenStudy (anonymous):

take all these probabilities and multiply them together to the get the probability of all three being red

OpenStudy (lexy148):

oh yeah thats right

OpenStudy (anonymous):

i.e. \[\frac{5}{10}\times \frac{4}{9}\times \frac{3}{8}\] cancel first, multiply last

OpenStudy (lexy148):

60/720

OpenStudy (lexy148):

wait im confused what do you cancel and multiply last

OpenStudy (anonymous):

you already said \(\frac{5}{10}=\frac{1}{2}\) right?

OpenStudy (lexy148):

yes

OpenStudy (anonymous):

just got logged out

OpenStudy (anonymous):

\[\frac{5}{10}\times \frac{4}{9}\times \frac{3}{8}\] every number in the numerator cancels with some number in the denominator your numerator will be 1 when you are finished cancelling hence "cancel first, multiply last"

OpenStudy (lexy148):

yes i got 1 for the numerator and 12 for the denominator

OpenStudy (anonymous):

good

OpenStudy (lexy148):

so would my answer be 1/12

OpenStudy (anonymous):

yes

OpenStudy (lexy148):

okay

OpenStudy (anonymous):

others work exactly the same way

OpenStudy (lexy148):

can you help me on the last one

OpenStudy (lexy148):

i find the last one more difficult

OpenStudy (anonymous):

no not at all c. the first is red, the second is white and the third is blue

OpenStudy (anonymous):

what is the probability the first is red?

OpenStudy (lexy148):

4/9

OpenStudy (anonymous):

?

OpenStudy (lexy148):

so 4/10 since they're asking for one red

OpenStudy (anonymous):

i think you may be confused by the question we are starting over

OpenStudy (anonymous):

A bag contains five red chips, three white chips, and two blue chips. Three chips are to be selected at random, without replacement. Find the probability that the first is red, the second is white and the third is blue

OpenStudy (anonymous):

so once again, what is the probability that the first one is red?

OpenStudy (lexy148):

1/10

OpenStudy (anonymous):

hmm no same as last time

OpenStudy (anonymous):

five out of ten are red

OpenStudy (lexy148):

ok i got that part

OpenStudy (anonymous):

now one red one is gone what is the probability that the next one is white?

OpenStudy (lexy148):

3/9

OpenStudy (anonymous):

yup one more and you are done

OpenStudy (anonymous):

one red one is gone one white one is gone what is the probability that the next one is blue?

OpenStudy (lexy148):

the next one would be 2/8

OpenStudy (anonymous):

so it would final job is \[\frac{5}{10}\times \frac{3}{9}\times \frac{2}{8}\] cancel first, multiply last

OpenStudy (lexy148):

ok i got 1/24 for my final answer